The intensity of a siren is 10.9 W/m2 when you are standing 20 m away from it. Now you go 120 m away. (in other words, 6 times farther away) How loud is it now?
Answer #1

this requires the use of the inverse square law;

sphere areaintensity at 4rtr2 surface of sphere source power 4str 4 A /A The energy twice as far from the 2r source is sprea

so, now,

at, r=20 m,I=10.9

P/(4*pi*r^2)= I

=>P=54789.37 watts

so, at 120 m

I=P/(4*pi*r^2)=0.302 W/m^2